OpenStudy (anonymous):

the table below lists recommended amounts of food to order for 25 party guests. How much of each food item should Nathan and Amanda order for a graduation party with 45 guests? use the table below ( item/amount) Friend chicken/40 pieces Lasagna / 14 pounds Deli meats / 4.5 pounds Sliced cheese / 2 1/3 Bakery Buns / 3 dozen Potato salad / 6 pounds So Nathan and Amanda should order ____ pieces of chicken, ___ pounds of lasagna,___ pounds of Deli meats,____ pounds of cheese,____ Dozen buns and ____ pounds of Potato salad?

OpenStudy (anonymous):

If 40 pieces feed 25 guests, then x pieces feed 45. Make a proportion... 40/25 = x/45 (cross multiply) 25x = (40)(45) 25x = 1800 x = 1800/25 x = 72 pieces of chicken. If 14 lbs of lasagna feed 25 people, then x lbs feed 45. 14/25 = x/45 25x = (14)(45) 25x = 630 x = 630/25 x = 25.2 lbs if 4.5 lbs of deli meat feet 25 people, then x lbs feed 45 4.5/25 = x/45 25x = (4.5)(45) 25x = 202.5 x = 202.5/25 x = 8.1 lbs. If 2.33 sliced cheese feeds 25 people, then x amount feeds 45 2.33/25 = x/45 25x = (2.33)(45) 25x = 104.85 x = 104.85/25 x = 4.19 lbs of sliced cheese If 3 doz buns feed 25 people, then x buns feeds 45 3/25 = x/45 25x = (3)(45) 25x = 135 x = 135/25 x = 5.4 doz bakery buns If 6 lbs of potato salad feed 25 people, then x lbs feed 45 people 6/25 = x/45 25x = (6)(45) 25x = 270 x = 270/25 x = 10.8 lbs of potato salad Nathan and Amanda should order 72 pieces of chicken, 25.2 pounds of lasagna, 8.1 pounds of deli meat, 4.19 pounds of sliced cheese, 5.4 dozen bakery buns, and 10.8 pounds of potato salad.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob. I hope they are correct :)